# Millen Kanabar and Michael Gastpar School of Computer and Communication Sciences, EPFL

### v3

@@ Appendix C Proofs of the oracle-assisted bounds
Denote the conditional distribution of each sample at stage $i$ given the auxiliary information $W_{i}^{n_{i}}$ as $\tilde{P}_{X|W}(p,\hat{P}_{i-1},\alpha)$, and for each $i\in[1:n]$, denote the joint distribution of $(X_{0}^{n_{0}},W_{1}^{n_{1}},X_{1}^{n_{1}},\dots,W_{i}^{n_{i}},X_{i}^{n_{i}})$ as $\bar{\mathbf{P}}^{\prime}_{i,(p)}$. Now for any $p_{1},p_{2}\in\Delta_{k}$, for every $i\in[0:t-1]$,
$D\left(\tilde{P}_{X|W}({p_{v},\hat{P}_{i},\alpha_{i+1}})\middle\|\tilde{P}_{X|W}({p_{v-2e_{j}},\hat{P}_{i},\alpha_{i+1}})\middle|P_{W_{i+1}^{n_{i+1}}},\bar{\mathbf{P}}^{\prime}_{i,(p_{v})}\right)=\alpha D(p_{v}\|p_{v-2e_{j}}).$ (23)
With the same construction of $\mathcal{P}_{k}$ as Section A, with the chain rule of the KL divergence along with (23), we find that
$D\Big{(}\bar{\mathbf{P}}^{\prime}_{t,(p_{v})}\Big{\|}\bar{\mathbf{P}}^{\prime}_{t,(p_{v-2e_{j}})}\Big{)}=$ $n_{0}D\left(p_{v}\|p_{v-2e_{j}}\right)+\sum_{i=0}^{t-1}n_{i+1}D\left(\tilde{P}_{X|W}({p_{v},\hat{P}_{i},\alpha_{i+1}})\middle\|\tilde{P}_{X|W}({p_{v-2e_{j}},\hat{P}_{i},\alpha_{i+1}})\middle|P_{W_{i+1}^{n_{i+1}}},\bar{\mathbf{P}}^{\prime}_{i,(p_{v})}\right)$
$=$ $D\left(p_{v}\|p_{v-2e_{j}}\right)\left(\sum_{i=0}^{t}n_{i}\alpha_{i}\right)\leq\left(\sum_{i=0}^{t}n_{i}\alpha_{i}\right)\frac{8\delta^{2}}{k(1-\delta^{2})}.$
@@ Appendix D Proof of the Lower Bound for the Non-Matching Regime
$\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}\leq$ $C\sum_{i=0}^{t}\frac{n_{i}\alpha_{i}}{\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}}$
$\leq$ $C_{\epsilon}+\epsilon C\sum_{i=0}^{t}\frac{n_{i}\alpha_{i}}{\sqrt{\sum_{j=0}^{i-1}n_{i}\alpha_{i}}}$
$\overset{(a)}{\leq}$ $C_{\epsilon}+\epsilon CD\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}},$
@@
$C_{\epsilon}=$ $\sum_{i=0}^{t_{\epsilon}}\frac{n_{i}\alpha_{i}}{\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}}$
and $\text{and }D=$ $\left(1+\sqrt{1+\sup_{s\geq 1}\frac{n_{s}\alpha_{s}}{\sum_{i=0}^{s-1}n_{i}\alpha_{i}}}\right).$
@@ Appendix E Proofs of Auxiliary Lemmas
We first prove the lemma for the $\ell_{2}^{2}$ loss. Define the sign()($\cdot$)( ⋅ ) function as
@@
and let ${.}\mathds{1}\{.\}$ be the indicator function.
@@
$r^{l_{2}^{2}}_{\lambda,n_{0},n_{1}}\geq$ $\frac{\delta^{2}}{k^{2}}\sum_{j=1}^{k/2}\frac{1}{2}\Bigg{[}\bar{\mathbf{Q}}_{t,(+j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=-1\right)$
$\qquad\qquad+\bar{\mathbf{Q}}_{t,(-j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=+1\right)\Bigg{]}$
$=$ $\frac{\delta^{2}}{2k^{2}}\sum_{j=1}^{k/2}\Bigg{[}1-\bar{\mathbf{Q}}_{t,(+j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=+1\right)$
$\qquad\qquad+\bar{\mathbf{Q}}_{t,(-j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=+1\right)\Bigg{]}$
@@
$E\left[\hat{P}_{0}[j]\hat{P}_{1,(\mathrm{cond})}[j]\right]=$ $E\left[E\left[\hat{P}_{0}[j]\hat{P}_{1,(\mathrm{cond})}[j]\middle|\hat{P}_{0}[j]\right]\right]$
$=$ $E\left[\frac{\hat{P}_{0}[j]}{\alpha}\left(\alpha p[j]+(1-\alpha)\hat{P}_{0}[j]-(1-\alpha)\hat{P}_{0}[j]\right)\right]$
@@
$=$ $E\left[\left(\alpha p[j]+(1-\alpha)\hat{P}_{0}[j]\right)^{2}\left(1-\frac{1}{n_{1}}\right)+\frac{1}{n_{1}}\left(\alpha p[j]+(1-\alpha)\hat{P}_{0}[j]\right)\right]-p[j]^{2}$
@@
$E\left[\hat{P}_{\mathrm{emp}}[j]\hat{P}_{0}[j]\right]-p[j]^{2}=$ $E\left[E\left[\hat{P}_{\mathrm{emp}}[j]\hat{P}_{0}[j]\big{|}\hat{P}_{0}[j]\right]\right]-p[j]^{2}$
@@
$E\left[\left(\hat{P}_{1,(\mathrm{cond})}[j]-p[j]\right)^{2}\right]=$ $\frac{1}{\alpha^{2}}\left(\text{var}\left(\hat{P}_{\mathrm{emp}}[j]\right)+(1-\alpha)^{2}\text{var}\left(\hat{P}_{0}[j]\right)-2(1-\alpha)\text{cov}\left(\hat{P}_{0}[j],\hat{P}_{\mathrm{emp}}[j]\right)\right)$
$=$ $\frac{p[j]\left(1-p[j]\right)}{\alpha^{2}}\left(\frac{1}{n_{1}}+\eta(1-\alpha)^{2}\left(1-\frac{1}{n_{1}}\right)+(1-\alpha)^{2}\cdot\eta-2\eta(1-\alpha)^{2}\right)$
@@
We proceed by induction. It is easy to verify (26) for $t=1$. Assume that it holds for $t=u-1$. Then
$\sum_{i=1}^{u}\frac{a_{i}}{\sqrt{\sum_{j=0}^{i-1}a_{j}}}\leq$ $D\sqrt{\sum_{i=0}^{u-1}a_{i}}+\frac{a_{u}}{\sqrt{\sum_{i=0}^{u-1}a_{i}}}=\frac{D\sum_{i=0}^{u-1}a_{i}+a_{u}}{\sqrt{\sum_{i=0}^{u-1}a_{i}}}.$

### v4

@@ Appendix C Proofs of the oracle-assisted bounds
Denote the conditional distribution of each sample at stage $i$ given the auxiliary information $W_{i}^{n_{i}}$ as $\tilde{P}_{X|W}(p,\hat{P}_{i-1},\alpha)$, and for each $i\in[1:n]$, denote the joint distribution of $(X_{0}^{n_{0}},W_{1}^{n_{1}},X_{1}^{n_{1}},\dots,W_{i}^{n_{i}},X_{i}^{n_{i}})$ as $\bar{\mathbf{P}}^{\prime}_{i,(p)}$. Now for any $p_{1},p_{2}\in\Delta_{k}$, for every $i\in[1:t]$,
$E\left[D\left(\tilde{P}_{X|W}({p_{v},\hat{P}_{i{-}1},\alpha_{i}})\middle\|\tilde{P}_{X|W}({p_{v{-}2e_{j}},\hat{P}_{i{-}1},\alpha_{i}})\right)\right]=\alpha D(p_{v}\|p_{v{-}2e_{j}}).$ (24)
where the expectation is taken with respect to both $P_{W_{i}^{n_{i}}}$ and $\bar{\mathbf{P}}^{\prime}_{i{-}1,(p_{v})}$. With the same construction of $\mathcal{P}_{k}$ as Section A, with the chain rule of the KL divergence along with (24), we find that
$D\Big(\bar{\mathbf{P}}^{\prime}_{t,(p_{v})}\Big\|\bar{\mathbf{P}}^{\prime}_{t,(p_{v-2e_{j}})}\Big)=$ $D\left(p_{v}\|p_{v-2e_{j}}\right)\left(\sum_{i=0}^{t}n_{i}\alpha_{i}\right)\leq\left(\sum_{i=0}^{t}n_{i}\alpha_{i}\right)\frac{8\delta^{2}}{k(1-\delta^{2})}.$
@@ Appendix D Proof of the Lower Bound for the Non-Matching Regime
$\sum_{i=0}^{t}n_{i}\alpha_{i}^{2}\leq$ $C\sum_{i=0}^{t}\frac{n_{i}\alpha_{i}}{\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}}\leq C_{\epsilon}+\epsilon C\sum_{i=0}^{t}\frac{n_{i}\alpha_{i}}{\sqrt{\sum_{j=0}^{i-1}n_{i}\alpha_{i}}}\overset{(a)}{\leq}C_{\epsilon}+\epsilon CD\sqrt{\sum_{i=0}^{t}n_{i}\alpha_{i}},$
@@
$C_{\epsilon}=$ $\sum_{i=0}^{t_{\epsilon}}\frac{n_{i}\alpha_{i}}{\sum_{j=0}^{i-1}n_{j}\alpha_{j}^{2}}\text{ and }D=\left(1+\sqrt{1+\sup_{s\geq 1}\frac{n_{s}\alpha_{s}}{\sum_{i=0}^{s-1}n_{i}\alpha_{i}}}\right).$
@@ Appendix E Proofs of Auxiliary Lemmas
We first prove the lemma for the $\ell_{2}^{2}$ loss. Define the sign()($\cdot$) function as
@@
and let $\mathds{1}\{.\}$ be the indicator function.
@@
$r^{l_{2}^{2}}_{\lambda,n_{0},n_{1}}\geq$ $\frac{\delta^{2}}{k^{2}}\sum_{j=1}^{k/2}\frac{1}{2}\Bigg[\bar{\mathbf{Q}}_{t,(+j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=-1\right)+\bar{\mathbf{Q}}_{t,(-j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=+1\right)\Bigg]$
$=$ $\frac{\delta^{2}}{2k^{2}}\sum_{j=1}^{k/2}\Bigg[1-\bar{\mathbf{Q}}_{t,(+j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=+1\right)+\bar{\mathbf{Q}}_{t,(-j)}\left(\mathrm{sign}\left(\hat{P}_{t}-\frac{1}{k}\right)=+1\right)\Bigg]$
@@
$E$ $\left[\hat{P}_{0}[j]\hat{P}_{1,(\mathrm{cond})}[j]\right]$
$=$ $E\left[E\left[\hat{P}_{0}[j]\hat{P}_{1,(\mathrm{cond})}[j]\middle|\hat{P}_{0}[j]\right]\right]$
$=$ $E\Bigg[\frac{\hat{P}_{0}[j]}{\alpha}\Big(\alpha p[j]+(1-\alpha)\hat{P}_{0}[j]-(1-\alpha)\hat{P}_{0}[j]\Big)\Bigg]$
@@
$=$ $E\Bigg[\left(\alpha p[j]+(1-\alpha)\hat{P}_{0}[j]\right)^{2}\left(1-\frac{1}{n_{1}}\right)+\frac{1}{n_{1}}\left(\alpha p[j]+(1-\alpha)\hat{P}_{0}[j]\right)\Bigg]-p[j]^{2}$
@@
$E\Big[\hat{P}_{\mathrm{emp}}[j]\hat{P}_{0}[j]\Big]-p[j]^{2}=$ $E\left[E\left[\hat{P}_{\mathrm{emp}}[j]\hat{P}_{0}[j]\big|\hat{P}_{0}[j]\right]\right]-p[j]^{2}$
@@
$E\Big[\left(\hat{P}_{1,(\mathrm{cond})}[j]-p[j]\right)^{2}\Big]=$ $\frac{1}{\alpha^{2}}\Big(\text{var}\left(\hat{P}_{\mathrm{emp}}[j]\right)+(1-\alpha)^{2}\text{var}\left(\hat{P}_{0}[j]\right)-2(1-\alpha)\text{cov}\left(\hat{P}_{0}[j],\hat{P}_{\mathrm{emp}}[j]\right)\Big)$
$=$ $\frac{p[j]\left(1-p[j]\right)}{\alpha^{2}}\Big(\frac{1}{n_{1}}+\eta(1-\alpha)^{2}\left(1-\frac{1}{n_{1}}\right)+(1-\alpha)^{2}\cdot\eta-2\eta(1-\alpha)^{2}\Big)$
@@
We proceed by induction. It is easy to verify (27) for $t=1$. Assume that it holds for t=u1t=u-1. Then
$\sum_{i=1}^{u}\frac{a_{i}}{\sqrt{\sum_{j=0}^{i-1}a_{j}}}\leq$ $D\sqrt{\sum_{i=0}^{u-1}a_{i}}+\frac{a_{u}}{\sqrt{\sum_{i=0}^{u-1}a_{i}}}$
$=$ $\frac{D\sum_{i=0}^{u-1}a_{i}+a_{u}}{\sqrt{\sum_{i=0}^{u-1}a_{i}}}.$